An equiconvex lens of RI 1.6 has power 4 D in air.Its power in water is ?

Grade:11

1 Answers

Arun
6 years ago
Dear Ayusha
 
RI = 1.6
P = 4D
From lens maker formula
P = (\mu-1) [1/R1 -  1/R2]
4 = (1.6-1) [1/R1  - 1/R2]
[1/R1 - 1/R2] = 4/0.6 = 20/3
 
RI of water  = 1.33
 
Refractive index of glass wrt to water = 1.6/1.33
 
Power of lens in water = (1.6/1.33  -1)(20/3)
= 1.3533 D 
 
Regards
Arun (askIITians forum expert)